What is the final temperature of 8575 cal of heat when removed from 15.0 g steam at 1000C?


1
Expert's answer
2021-03-21T11:31:13-0400

"\Delta{Q}= 8575 \ cal;\Delta{Q}\approx4.2*8575=36015\ J=36\ kJ"

"m= 15g = 0.015\ kg"

"T = 100\degree +273=373\ K"

"Q_1 \text{ the amount of heat in the transition from steam to liquid}"

"L=2260\frac{kJ}{kg}"

"Q_1=Lm= 2260*0.015=33.9\ kJ"

"Q_2= \Delta{Q}-Q_1=36-33.9= 2.1\ kJ"

"Q_2\text{ the amount of heat when the temperature changes}"

"Q_2= cm\Delta{T}"

"c= 4.2\frac{kJ}{kg*K}"

"\Delta{T}= \frac{Q_2}{cm}=\frac{2.1}{4.2*0.015}=33\ K"

"T_2= T-\Delta{T}=373-33= 340\ K=67\degree"

Answer: "67\degree\text{ final temperature}"